OpenStudy (anonymous):

What is the solution to the rational equation 3 over 2 - 1 over x plus 5 = negative 2 over 2x plus 10 ?

OpenStudy (doc.brown):

Let me just write it out first, \[\frac{3}{2}-\frac{1}{x+5}=\frac{-2}{2x+10}\]

OpenStudy (doc.brown):

Could 2x+10 be a common denominator?

OpenStudy (anonymous):

@doc.brown These are the options given: x = -5 x = -5 over 2 x = 5 No solution

OpenStudy (doc.brown):

This is how I did it \[\frac{x+10}{x+10}(\frac{3}{2})-\frac{2}{2}(\frac{1}{x+5})=\frac{-2}{2x+10}\]\[\frac{3x+10}{2x+10}-\frac{2}{2x+10}=\frac{-2}{2x+10}\]

OpenStudy (anonymous):

@doc.brown So out of those four options which one do you think it would be then? Still kind of confused on how to get the answer?

OpenStudy (doc.brown):

Ok, watch what happens when you add 2/(2x+10) to both sides\[\frac{3x+10}{2x+10}-\frac{2}{x+10} + \frac{2}{x+10}=\frac{-2}{2x+10} +\frac{2}{x+10}\]\[\frac{3x+10}{2x+10}=0\]